OpenStudy (anonymous):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work.

OpenStudy (anonymous):

I have already gotten the equation but now I need to solve it and I cant figure it out. the equation is: 75/100*x > 25/100*x

OpenStudy (***[isuru]***):

\[\frac{ the \ volume \ of \ apple \ juice }{ the \ total \ volume \ } \times 100 = 75\] let \[ the \ volume \ apple \ juice\ = x\] then \[\frac{ x }{ 1+ x } \times 100 = 75\] and the expression is 100x = 75 ( 1+ x) got it ?

OpenStudy (anonymous):

yea kinda. would the 100x=75(1+x) be what I need to solve.?

OpenStudy (***[isuru]***):

yes...

OpenStudy (anonymous):

ok so it would be 100x=75(1+x) 100x=75+75 100x+75=75+75x+75 175x=75

OpenStudy (anonymous):

then I divide both sides by 75.

OpenStudy (***[isuru]***):

no... wait

OpenStudy (***[isuru]***):

u have made a mistake.. check ur steps again

OpenStudy (anonymous):

I don't see where it is.....

OpenStudy (anonymous):

the first thing I would have to do is take 75 and multiply it to whats in the parenthesis which is gonna make it where the equation is 100x=75+75x

OpenStudy (***[isuru]***):

100x = 75 + 75 x 100x -75x = 75 + 75x -75x 25x = 75 got it ?

OpenStudy (anonymous):

yes..

OpenStudy (***[isuru]***):

then u will get x = 3 liters... tada ..... problem SOLVED XD

OpenStudy (anonymous):

lol (: thank you.! (;